OpenStudy (anonymous):

Find the root of this equation: log(x-2)=1

Parth (parthkohli):

Base \(10\)? You will have:\[\log(x - 2) = 1\iff 10^1 = x - 2\]I'm sure it is easy enough for you to follow.

OpenStudy (anonymous):

Here's an example of the method my teacher would like for me to follow

OpenStudy (anonymous):

x=12 ?

Parth (parthkohli):

Have you heard about the exponential form to logarithmic form? Basically, you can show this:\[\log(x - 2) = 1 \tag{1}\]\(\textbf{AXIOM:}\) \(\log_a b = c \iff a^c = b\).\[\therefore \log (x - 2) = 1 \iff 10^1 = x-2\]

Parth (parthkohli):

Yes, \(x = 12\).

OpenStudy (anonymous):

Oh yes thank you:) I was just confused because I usually have question involving the log of both sides

Parth (parthkohli):

Oh, good enough. :-)

OpenStudy (anonymous):

@ParthKohli how would i find the restrictions ?

Parth (parthkohli):

Restrictions? For the function \(f(x) = \log(x - 2)\)?

OpenStudy (anonymous):

Yes

Parth (parthkohli):

\(\log(x)\) is restricted to non-negative values for \(x\).

Parth (parthkohli):

So when \(x - 2 \ge 0\), the function is defined.

OpenStudy (anonymous):

look in the attachment i posted earlier

OpenStudy (anonymous):

Kind of like that

Parth (parthkohli):

Actually the restrictions are \(x < 0\) instead of \(x>0\).

OpenStudy (agent0smith):

@parthkohli "So when x−2≥0, the function is defined." I think that should be when x-2>0, as log x is only defined for x>0.

Parth (parthkohli):

@agent0smith Hmm, that's a very active argument. But,\[\log(0) = -\infty\]